A firm produces 40,000 pieces per month. The company has decided to assess the emerging demand...

  1. A firm produces 40,000 pieces per month. The company has decided to assess the emerging demand for the items in the next five years. The company forecast reveals that there is 40% probability that the demand will be strong during the planning period and 60% probability that the demand will be moderate. The company can add new capacity to meet the demand or expand the existing factory or to go for subcontracting and if the company has a strong growth in demand, then the new capacity could be added a year later. The cost of adding new capacity is Rs. 7,50,000 and this cost goes up by 5% if it is deferred by a year. The cost of expanding in the existing factory itself is Rs.2,75,000. The cost of subcontracting is negligible. The revenue accruing for each option is as follow,
  1. From new capacity: If the growth is strong, the revenue will be Rs.8,50,000 and in the case of a moderate growth, the revenue will be Rs. 4,00,000. These figures do not change even if the new unit comes into operation a year later.
  2. From expansion of the existing capacity: If the growth is strong, the revenue will be Rs.5,50,000 and in the case of a moderate growth, the revenue will be Rs. 3,00,000.
  3. From existing factory with subcontracting: If the growth is strong, the revenue will be Rs.3,50,000 and in the case of a moderate growth, the revenue will be Rs. 1,80,000.

Construct a decision tree to decide on the best capacity planning strategy. Report the same.
Answer #1

The decision tree is as follows:

Strong growth 850,000 100,000 0.00 0.40 Add new capacity -750,000 0.00 580,000 Moderate growth -350,000 0.00 400,000 0.60 Str

Expected Monetary Value (EMV) of each decision are computed as under:

EMV of Add new capacity = 850000*0.4+400000*0.6-750000 = 580000-750000 = Rs -1,70,000  (loss)

-

EMV of Expand existing facility = 550000*0.4+300000*0.6-275000 = 400000-275000 = Rs 1,25,000

-
Payoff, in case of Existing factory with Subcontracting and Add new capacity after one year, if strong growth = 850000 - 750000*1.05 = Rs 62,500

Payoff Don't add new capacity is higher (350,000) . Therefore, if growth is strong, then optimum decision is Don't add new capacity.

EMV of Existing factory with subcontracting = 0.4*350000+0.6*180000 = Rs 2,48,000

-

EMV of Existing factory with subcontracting is the highest of all decisions.

Therefore, optimum decision is Existing factory with subcontract, and if growth is strong, then Don't add new capacity

EMV of this decision = Rs 2,48,000